[Brahms-l] Shift report 20050131 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Jan 31 2005 - 16:26:31 EST
Had beam until noon.  The store was aborted and then MCR took over the machine for machine development.  FS at 8 deg A1219 and MRS 45 deg B2225
Runs achieved:
Run   tr2     tr3
12711 40k    140k
12712 35k    254k
12713 22k    159k
Then 2 min bias runs
Run   tr1   tr4   tr5
12714  304333  343789  281390
12717  96089   96337    74063

Dana moved the MRS to 40 deg after above was finished
I changed the MRS field to B1050


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jan 31 16:26:58 2005

This archive was generated by hypermail 2.1.8 : Mon Jan 31 2005 - 16:27:25 EST